# Mitigating Forgetting in LLM Fine-Tuning via Low-Perplexity Token Learning

### v6

@@ 1 Introduction
Supervised fine-tuning large language models (LLMs) [29] has proven highly effective for enhancing their ability to follow novel instructions [43, 21, 23, 15, 49] and produce useful outputs across a wide range of tasks, including summarization and web querying [34, 26, 1, 17]. As LLMs are increasingly applied in specialized domains such as arithmetic and programming assistance [38, 27], practitioners often encounter substantial computational and data cleaning issues when adapting these models to gain improvements. These difficulties are exacerbated by performance saturation in smaller LLMs, where benchmark improvements on tasks like MATH [12] and MBPP [3] tend to plateau as model size decreases. Moreover, fine-tuning can risk degrading a model’s general capabilities [31, 24, 30, 4]. This raises a critical question: how can we efficiently fine-tune pre-trained, instruction-following LLMs for domain-specific tasks without compromising other skill sets such as arithmetic or common sense reasoning, especially when the original fine-tuning data is unavailable or computational resources are constrained?
Recent work has shown that fine-tuning with model-generated data can be highly effective, often outperforming training on ground truth data when using high-quality generated samples [48, 10, 32, 35]. For example, Ren et al. [31] demonstrated that model-generated responses can surpass human annotations in downstream performance, particularly when using more capable models (e.g., GPT-4) to generate training data and mitigate degradation on general tasks. Although they attributed part of this success to the “familiarity” of generated responses, measured via low perplexity, it remains unclear why such data works well for instruction tuning. Specifically: (1) The regenerated response may differ contextually from the ground truth, and such contextual bias could affect the utility of low-perplexity training. (2) The “non-forgetting” impact of low-perplexity training on the performance of tasks not being fine-tuned, which maintains the practicality of the model, is still underexplored. (3) Since model size plays a role in distillation benefits, it is important to disentangle the effects of model scale and evaluate low-perplexity training using models of the same size.
@@ 2.1 Training and evaluation framework
For target-task fine-tuning, we adopt the MBPP and MATH datasets. These datasets provide rich annotations, including full solutions, reasoning steps, and test cases, beyond just the final answers, allowing models to learn comprehensive task-solving behaviors. For evaluation, in addition to using the test sets from MBPP and MATH, we assess model performance on GSM8K [7], ARC-Challenge [6], and BIRD [19] to examine generalization to non-target tasks involving various forms of reasoning and generation. A detailed description of the datasets is provided in Section 4.
@@ 2.2 Self-Output
Self-Output follows a similar high-level approach, similar to other synthetic data generation techniques [48, 31], where a high-quality synthetic dataset is created by sampling diverse responses from an LLM and applying a strict filtering process. For each training instance, we use a language model $M$ (e.g., Llama 3 8B Instruct) to generate $N$ (e.g., 32) distinct responses with a temperature setting of $T=0.7$. The generated responses are then filtered to those that are semantically aligned with the ground truth, ensuring the quality and relevance of synthetic data.
@@ 2.3 Rephrase
Yang et al. [47] proposed self-distillation. The process uses an instruction-finetuned LLM to rephrase ground-truth responses in its style. By providing both the instruction and ground truth to the target LLM, it generates semantically equivalent reformulations with ground truth responses. Compared to Self-Output, Rephrase is computationally efficient, as it requires only a single pass of generation without additional filtering. However, Rephrase is susceptible to hallucinations, as evidenced by the higher variance in token-level perplexity (see Table  1). Despite this limitation, Rephrase offers versatility, as it does not rely on verifiable ground truth and can be applied to a broader range of tasks.
@@
Table 2 presents the performance comparison between Self-Output, Rephrase, and Baseline Fine-tuning on ground truth using different models in terms of model size and model series across both target task improvement and non-target task degradation rate. Both task improvement rate (TI) and degradation rate (BWT) are adopted from Backward Transfer rate from [14] and modified as follows:
@@
Given $T$ tasks composed of one target task and T1T-1 non-target tasks, TI is the target task performance improvement percentage after training on the target task. BWT refers to the average performance degradation percentage of the T1T-1 non-target task after training on the target task. The models to fine-tune in this paper refer to generic instruction following models before fine-tuning with target task data, e.g., Llama 3 8B Instruct [8], Gemma 2 IT 2B [36], Mistral 7B Instruct [16], and OLMo 2 7B series [40]. For target task performance, Self-Output demonstrates consistent positive TI in both MBPP and MATH. For the non-target task, the Self-Output method achieves better BWT scores than the baseline fine-tuning in most cases, indicating much less degradation on non-target task capabilities. For more details, such as performance comparison with other models or other minor findings for self-output and rephrase method, please refer to Appendix D.
@@ 3 Low-perplexity token learning via selective token masking
While our approach shares conceptual similarities with prior work such as [20, 2], it differs significantly in both implementation and efficiency. Previous methods typically involved a two-stage process: training a reference model on high-quality data to identify learnable (high-perplexity) tokens, followed by pre-training a larger LLM on this curated subset. In contrast, STM achieves similar goals in a streamlined single-stage workflow by directly using an existing instruction-tuned model to compute token perplexities in a single forward pass. This design avoids the need to train or maintain any additional models, significantly reducing computational overhead while preserving the benefits of low-perplexity training. This makes STM a compelling choice for efficient fine-tuning rather than full-scale pretraining.
@@ 4 Experimental setting
To evaluate the effectiveness of our proposed method, we adopt benchmark datasets specifically selected to assess model robustness on both target and non-target tasks after fine-tuning. We include only datasets with verifiable ground truth, allowing us to filter out incorrect self-generated responses for a fair comparison with self-synthesized baselines and to ensure high data quality. Following recent findings [4] that LoRA [13] effectively mitigates performance degradation, we standardize our experiments by using LoRA for all fine-tuning. We then vary training strategies and data sources to study their impact on performance preservation. Details of our evaluation prompts and training configurations are provided in Appendix F. All implementation code and datasets are publicly available at: https://github.com/appier-research/robust-llm-finetunes.
@@ 5.4 Is forgetting affected by smaller learning rates?
Following training settings in [9, 37], learning rate (lr) would be one reason that harms the performance of Baseline Fine-tuning severely (but not for STM and LLM generated data baselines). Therefore, we conducted additional learning rate sweeps down to 1e71e-7 for Llama 3 8B Instruct and Gemma 2 IT 2B on the MBPP dataset. As Table 5 shows, we found that Baseline Fine-tuning improves at smaller learning rates (e.g., TI: $10.6\%$$1.33\%$), but the forgetting problem (BWT) persists (e.g., $1.6\%$). In contrast, STM consistently outperforms Baseline across all learning rates with stronger target task performance and substantially lower forgetting. Notably, STM achieves better BWT and stable performance even at small learning rates (e.g., TI: $2.23\%$, BWT: $1.39\%$ at 1e-7). These results demonstrate that STM is more robust to learning rate choices, requires less learning rate tuning, and achieves more stable training dynamics, highlighting its practical advantage in real-world fine-tuning scenarios.
@@ 5.5 STM generalization on different fine-tuning strategies
As STM is a masking technique applied to target data during the fine-tuning stage, it should be orthogonal to the fine-tuning techniques and improve the performance degradation. Thus, we experiment with how STM improves practical fine-tuning strategies like Full-Weight Fine-tuning (FWFT), Low-rank adaptation (LoRA), and Weight-decomposed LoRA (DoRA) [22]. As Table 4 shows, all of the fine-tuning techniques are further improved by STM, which indicates STM’s generalization on different fine-tuning strategies.
@@ 6 Analysis and discussion
Although we found that directly training with high perplexity tokens leads to higher loss and more weights updated. Applying regularization techniques can reduce model weight changes to mitigate non-target task degradation. Hence, we apply alternative regularization techniques to fine-tune with high-perplexity ground truth data to mitigate non-target task performance degradation. Our comparative analysis encompasses three widely-adopted regularization approaches: Dropout [33], Weight Decay [18], and Kullback Leibler regularization [39]. These methods have demonstrated effectiveness in preventing models from overfitting. Table 8 shows that under similar model weight changes, models applied with STM still outperform the high perplexity data trained models with regularization, indicating the performance degradation results from high perplexity training instead of purely model overfitting. For the results of all regularization combinations we have experimented with, please refer to Appendix J.
@@ 7 Related work
Instruction following [28, 5] is widely used to align responses of LLMs with target task values. Practically, it requires instruction tasks dataset to fine-tune LLMs, where each dataset consists of instructions and desired responses. Hence, the performance of instruction-tuning heavily relies on the quality of instruction data such as context richness [45, 41, 50]. In addition, Ghosh et al. [9] investigated the limitation of instruction-tuning of LLMs about the catastrophic forgetting from pattern-copying behaviors and hallucinations with Lora/full fine-tuning. These works give us a good start on how performance degrades in terms of models’ response behaviors and benchmarks when fine-tuning with instruction following dataset.
@@
As Wang et al. [41], Chung et al. [5] showed, LLMs break down easily after training with different tasks. Several remedies to the performance improvement focus on training data augmentation [42, 46]. For instance, [31] uses much larger LLMs (e.g., GPT-4 and Claude) to generate responses of questions as training labels, which improves the performance on both the target task and other non-target tasks. However, such a distillation method neglects the correctness of generated labels, so that more incorrect responses could be trained as the amount of generated data increases, and thus, using equal or smaller-sized models for distillation could be challenging. [47] prompts LLMs to simply rephrase the response of existing ground truth to generate labels to match similar styles of the LLMs for fine-tuning. However, rephrasing the ground truth answer limits the output distribution and results in lower performance in our study. Furthermore, [11] exploits a base LLM as a judge to pick out answerable and unanswerable questions to compose a new training dataset to improve target task performance only on the QA/conversation dataset. Although using Mistral 7B Instruct to generate acceptable responses for answerable questions improves target and non-target tasks, it is rarely discussed that if the proposed method is applicable to different model sizes and series. Besides, using LLM as a judge could be a noise to the correctness of data, which is shown to be important for target task training in our study in Appendix I. Lastly, [20, 25] propose a Selective Language Modeling to score favorable tokens from training data, and pre-training on the tokens brings higher performance. However, the requirements for pre-training a reference model or preparing a high-quality dataset for a scoring model could be exhausting depending on task difficulty. In addition to STM’s applicability in Fine-Tuning, STM does not require pre-training for the selection of tokens at all, which brings efficiency and performance improvement at the same time.
@@ Appendix B Exploration on additional tasks
For instruction following evaluation, we use IFEval dataset [51], which consist of 541 prompts from 25 verifiable instruction type, as an additional non-target task evaluation for MBPP trained models.
@@
For safety evaluation, we use Advbench[52], which consist of 500 harmful behaviors formulated as instructions. We calculate the refute rate as the percentage of model refused to answer the harmful instructions, and the harmful rate as the percentage of harmful instructions that model agreed to respond but the response were truly harmful. We use GPT4.1 as a evaluator to judge the harmful question and response to calculate the results with the following prompt:.
@@ F.3 Evaluation prompt
In BIRD task, we basically follow the same prompt as [44] does. A database schema (schema) is followed by a text-to-sql instruction and the question query (question) as the Textbox below shows:
@@ Appendix G Other STM alternatives
Besides cross-scale setting, inspired by [20, 25], we also investigate possible STM settings that acquire a model trained on ground truth data:
@@ Appendix H A continual learning setting of STM
We have experimented the effectiveness of STM by training on a single target task to reduce non-target tasks degradation. Yet, in a setting of continual learning scenario as [14] presents, degradation of tasks also occurs after several target tasks are trained sequentially. We simulate such setting to fine-tune Llama 3 8B Instruct with MATH initially and followed by MBPP tasks. Table 19 shows that (1) not only the target task is improved after fine-tuning, but also its degradation is very small after training another target task; (2) the non-target tasks like GSM8K, ARC and BIRD keeps similar performance even after two rounds of fine-tuning. It indicates STM can be effective in such continual learning setting as well.

### v7

@@ 1 Introduction
Supervised fine-tuning large language models (LLMs) [ouyang2022training] has proven highly effective for enhancing their ability to follow novel instructions [zsfewshot, liu2025synthesizing, lu2024discussion, hung2025simtube, zhang2023bootstrap] and produce useful outputs across a wide range of tasks, including summarization and web querying [stiennon2020learning, nakano2021webgpt, achiam2023gpt, kopf2024openassistant]. As LLMs are increasingly applied in specialized domains such as arithmetic and programming assistance [trummer2022codexdb, 10.1145/3524842.3528470], practitioners often encounter substantial computational and data cleaning issues when adapting these models to gain improvements. These difficulties are exacerbated by performance saturation in smaller LLMs, where benchmark improvements on tasks like MATH [Hendrycks2021MeasuringMP] and MBPP [austin2021program] tend to plateau as model size decreases. Moreover, fine-tuning can risk degrading a model’s general capabilities [ren2024learn, mai2024finetuningfinecalibrated, qi2023finetuningalignedlanguagemodels, biderman2024lora]. This raises a critical question: how can we efficiently fine-tune pre-trained, instruction-following LLMs for domain-specific tasks without compromising other skill sets such as arithmetic or common sense reasoning, especially when the original fine-tuning data is unavailable or computational resources are constrained?
Recent work has shown that fine-tuning with model-generated data can be highly effective, often outperforming training on ground truth data when using high-quality generated samples [zelikman2022star, Gulcehre2023ReinforcedS, Singh2023BeyondHD, suwono2023location]. For example, ren2024learn demonstrated that model-generated responses can surpass human annotations in downstream performance, particularly when using more capable models (e.g., GPT-4) to generate training data and mitigate degradation on general tasks. Although they attributed part of this success to the “familiarity” of generated responses, measured via low perplexity, it remains unclear why such data works well for instruction tuning. Specifically: (1) The regenerated response may differ contextually from the ground truth, and such contextual bias could affect the utility of low-perplexity training. (2) The “non-forgetting” impact of low-perplexity training on the performance of tasks not being fine-tuned, which maintains the practicality of the model, is still underexplored. (3) Since model size plays a role in distillation benefits, it is important to disentangle the effects of model scale and evaluate low-perplexity training using models of the same size.
@@ 2.1 Training and evaluation framework
For target-task fine-tuning, we adopt the MBPP and MATH datasets. These datasets provide rich annotations, including full solutions, reasoning steps, and test cases, beyond just the final answers, allowing models to learn comprehensive task-solving behaviors. For evaluation, in addition to using the test sets from MBPP and MATH, we assess model performance on GSM8K [cobbe2021gsm8k], ARC-Challenge [clark2018think], and BIRD [li2024can] to examine generalization to non-target tasks involving various forms of reasoning and generation. A detailed description of the datasets is provided in Section 4.
@@ 2.2 Self-Output
Self-Output follows a similar high-level approach, similar to other synthetic data generation techniques [zelikman2022star, ren2024learn], where a high-quality synthetic dataset is created by sampling diverse responses from an LLM and applying a strict filtering process. For each training instance, we use a language model $M$ (e.g., Llama 3 8B Instruct) to generate $N$ (e.g., 32) distinct responses with a temperature setting of $T=0.7$. The generated responses are then filtered to those that are semantically aligned with the ground truth, ensuring the quality and relevance of synthetic data.
@@ 2.3 Rephrase
Yang2024SelfDistillationBD proposed self-distillation. The process uses an instruction-finetuned LLM to rephrase ground-truth responses in its style. By providing both the instruction and ground truth to the target LLM, it generates semantically equivalent reformulations with ground truth responses. Compared to Self-Output, Rephrase is computationally efficient, as it requires only a single pass of generation without additional filtering. However, Rephrase is susceptible to hallucinations, as evidenced by the higher variance in token-level perplexity (see Table  1). Despite this limitation, Rephrase offers versatility, as it does not rely on verifiable ground truth and can be applied to a broader range of tasks.
@@
Table 2 presents the performance comparison between Self-Output, Rephrase, and Baseline Fine-tuning on ground truth using different models in terms of model size and model series across both target task improvement and non-target task degradation rate. Both task improvement rate (TI) and degradation rate (BWT) are adopted from Backward Transfer rate from [huang2024mitigatingcatastrophicforgettinglarge] and modified as follows:
@@
Given $T$ tasks composed of one target task and T1T-1 non-target tasks, TI is the target task performance improvement percentage after training on the target task. BWT refers to the average performance degradation percentage of the T1T-1 non-target task after training on the target task. The models to fine-tune in this paper refer to generic instruction following models before fine-tuning with target task data, e.g., Llama 3 8B Instruct [dubey2024llama], Gemma 2 IT 2B [team2024gemma], Mistral 7B Instruct [jiang2023mistral], and OLMo 2 7B series [olmo20252olmo2furious]. For target task performance, Self-Output demonstrates consistent positive TI in both MBPP and MATH. For the non-target task, the Self-Output method achieves better BWT scores than the baseline fine-tuning in most cases, indicating much less degradation on non-target task capabilities. For more details, such as performance comparison with other models or other minor findings for self-output and rephrase method, please refer to Appendix D.
@@ 3 Low-perplexity token learning via selective token masking
While our approach shares conceptual similarities with prior work such as [linnot, ankner2025perplexed], it differs significantly in both implementation and efficiency. Previous methods typically involved a two-stage process: training a reference model on high-quality data to identify learnable (high-perplexity) tokens, followed by pre-training a larger LLM on this curated subset. In contrast, STM achieves similar goals in a streamlined single-stage workflow by directly using an existing instruction-tuned model to compute token perplexities in a single forward pass. This design avoids the need to train or maintain any additional models, significantly reducing computational overhead while preserving the benefits of low-perplexity training. This makes STM a compelling choice for efficient fine-tuning rather than full-scale pretraining.
@@ 4 Experimental setting
To evaluate the effectiveness of our proposed method, we adopt benchmark datasets specifically selected to assess model robustness on both target and non-target tasks after fine-tuning. We include only datasets with verifiable ground truth, allowing us to filter out incorrect self-generated responses for a fair comparison with self-synthesized baselines and to ensure high data quality. Following recent findings [biderman2024lora] that LoRA [hu2022lora] effectively mitigates performance degradation, we standardize our experiments by using LoRA for all fine-tuning. We then vary training strategies and data sources to study their impact on performance preservation. Details of our evaluation prompts and training configurations are provided in Appendix F. All implementation code and datasets are publicly available at: https://github.com/appier-research/robust-llm-finetunes.
@@ 5.4 Is forgetting affected by smaller learning rates?
Following training settings in [ghosh2024a-acloserlook, touvron2023llama2openfoundation], learning rate (lr) would be one reason that harms the performance of Baseline Fine-tuning severely (but not for STM and LLM generated data baselines). Therefore, we conducted additional learning rate sweeps down to 1e71e-7 for Llama 3 8B Instruct and Gemma 2 IT 2B on the MBPP dataset. As Table 5 shows, we found that Baseline Fine-tuning improves at smaller learning rates (e.g., TI: $10.6\%$$1.33\%$), but the forgetting problem (BWT) persists (e.g., $1.6\%$). In contrast, STM consistently outperforms Baseline across all learning rates with stronger target task performance and substantially lower forgetting. Notably, STM achieves better BWT and stable performance even at small learning rates (e.g., TI: $2.23\%$, BWT: $1.39\%$ at 1e-7). These results demonstrate that STM is more robust to learning rate choices, requires less learning rate tuning, and achieves more stable training dynamics, highlighting its practical advantage in real-world fine-tuning scenarios.
@@ 5.5 STM generalization on different fine-tuning strategies
As STM is a masking technique applied to target data during the fine-tuning stage, it should be orthogonal to the fine-tuning techniques and improve the performance degradation. Thus, we experiment with how STM improves practical fine-tuning strategies like Full-Weight Fine-tuning (FWFT), Low-rank adaptation (LoRA), and Weight-decomposed LoRA (DoRA) [dora]. As Table 4 shows, all of the fine-tuning techniques are further improved by STM, which indicates STM’s generalization on different fine-tuning strategies.
@@ 6 Analysis and discussion
Although we found that directly training with high perplexity tokens leads to higher loss and more weights updated. Applying regularization techniques can reduce model weight changes to mitigate non-target task degradation. Hence, we apply alternative regularization techniques to fine-tune with high-perplexity ground truth data to mitigate non-target task performance degradation. Our comparative analysis encompasses three widely-adopted regularization approaches: Dropout [srivastava2014dropout], Weight Decay [krogh1991simple], and Kullback Leibler regularization [vieillard2020leverage]. These methods have demonstrated effectiveness in preventing models from overfitting. Table 8 shows that under similar model weight changes, models applied with STM still outperform the high perplexity data trained models with regularization, indicating the performance degradation results from high perplexity training instead of purely model overfitting. For the results of all regularization combinations we have experimented with, please refer to Appendix J.
@@ 7 Related work
Instruction following [Ouyang2022TrainingLM, Chung2022ScalingIL] is widely used to align responses of LLMs with target task values. Practically, it requires instruction tasks dataset to fine-tune LLMs, where each dataset consists of instructions and desired responses. Hence, the performance of instruction-tuning heavily relies on the quality of instruction data such as context richness [Xu2023WizardLMEL, Wang2023HowFC, zhou2023lima]. In addition, ghosh2024a-acloserlook investigated the limitation of instruction-tuning of LLMs about the catastrophic forgetting from pattern-copying behaviors and hallucinations with Lora/full fine-tuning. These works give us a good start on how performance degrades in terms of models’ response behaviors and benchmarks when fine-tuning with instruction following dataset.
@@
As Wang2023HowFC, Chung2022ScalingIL showed, LLMs break down easily after training with different tasks. Several remedies to the performance improvement focus on training data augmentation [wang-etal-2023-self-instruct, magpie]. For instance, [ren2024learn] uses much larger LLMs (e.g., GPT-4 and Claude) to generate responses of questions as training labels, which improves the performance on both the target task and other non-target tasks. However, such a distillation method neglects the correctness of generated labels, so that more incorrect responses could be trained as the amount of generated data increases, and thus, using equal or smaller-sized models for distillation could be challenging. [Yang2024SelfDistillationBD] prompts LLMs to simply rephrase the response of existing ground truth to generate labels to match similar styles of the LLMs for fine-tuning. However, rephrasing the ground truth answer limits the output distribution and results in lower performance in our study. Furthermore, [gupta2024selectiveselfrehearsalfinetuningapproach] exploits a base LLM as a judge to pick out answerable and unanswerable questions to compose a new training dataset to improve target task performance only on the QA/conversation dataset. Although using Mistral 7B Instruct to generate acceptable responses for answerable questions improves target and non-target tasks, it is rarely discussed that if the proposed method is applicable to different model sizes and series. Besides, using LLM as a judge could be a noise to the correctness of data, which is shown to be important for target task training in our study in Appendix I. Lastly, [linnot, mindermann2022prioritized] propose a Selective Language Modeling to score favorable tokens from training data, and pre-training on the tokens brings higher performance. However, the requirements for pre-training a reference model or preparing a high-quality dataset for a scoring model could be exhausting depending on task difficulty. In addition to STM’s applicability in Fine-Tuning, STM does not require pre-training for the selection of tokens at all, which brings efficiency and performance improvement at the same time.
@@ Appendix B Exploration on additional tasks
For instruction following evaluation, we use IFEval dataset [zhou2023instructionfollowingevaluationlargelanguage], which consist of 541 prompts from 25 verifiable instruction type, as an additional non-target task evaluation for MBPP trained models.
@@
For safety evaluation, we use Advbench[zou2023universal], which consist of 500 harmful behaviors formulated as instructions. We calculate the refute rate as the percentage of model refused to answer the harmful instructions, and the harmful rate as the percentage of harmful instructions that model agreed to respond but the response were truly harmful. We use GPT4.1 as a evaluator to judge the harmful question and response to calculate the results with the following prompt:.
@@ F.3 Evaluation prompt
In BIRD task, we basically follow the same prompt as [wu2024streambenchbenchmarkingcontinuousimprovement] does. A database schema (schema) is followed by a text-to-sql instruction and the question query (question) as the Textbox below shows:
@@ Appendix G Other STM alternatives
Besides cross-scale setting, inspired by [linnot, mindermann2022prioritized], we also investigate possible STM settings that acquire a model trained on ground truth data:
@@ Appendix H A continual learning setting of STM
We have experimented the effectiveness of STM by training on a single target task to reduce non-target tasks degradation. Yet, in a setting of continual learning scenario as [huang2024mitigatingcatastrophicforgettinglarge] presents, degradation of tasks also occurs after several target tasks are trained sequentially. We simulate such setting to fine-tune Llama 3 8B Instruct with MATH initially and followed by MBPP tasks. Table 19 shows that (1) not only the target task is improved after fine-tuning, but also its degradation is very small after training another target task; (2) the non-target tasks like GSM8K, ARC and BIRD keeps similar performance even after two rounds of fine-tuning. It indicates STM can be effective in such continual learning setting as well.